# HuRef: HUman-REadable Fingerprint for Large Language Models

### v1

@@ Abstract
Protecting the copyright of large language models (LLMs) has become crucial due to their resource-intensive training and accompanying carefully designed licenses. However, identifying the original base model of an LLM is challenging due to potential parameter alterations through fine-tuning or continued pretraining. In this study, we introduce HuRef, a human-readable fingerprint for LLMs that uniquely identifies the base model without exposing model parameters or interfering with training. We first observe that the vector direction of LLM parameters remains stable after the model has converged during pretraining, showing negligible perturbations through subsequent training steps, including continued pretraining, supervised fine-tuning (SFT), and RLHF, which makes it a sufficient condition to identify the base model. The necessity is validated by continuing to train an LLM with an extra term to drive away the model parameters’ direction and the model becomes damaged. However, this direction is vulnerable to simple attacks like dimension permutation or matrix rotation, which significantly change it without affecting performance. To address this, leveraging the Transformer structure, we systematically analyze potential attacks and define three invariant terms that identify an LLM’s base model. We make these invariant terms human-readable by mapping them to a Gaussian vector using a convolutional encoder and then converting it into a natural image with StyleGAN2. The encoder discriminates between invariants from different base models and ensures Gaussian output through adversarial training, while StyleGAN2 transforms Gaussian vectors into dog images. Consequently, our method generates a dog image as an identity fingerprint for an LLM, where the dog’s appearance strongly indicates the LLM’s base model. Specifically, if the LLM is adapted from another base model, the generated dog highly resembles that model; otherwise if trained independently from scratch, it exhibits a unique dog image distinct from other models. Experimental results across various LLMs demonstrate the effectiveness of our method, the generated dog image remains invariant to different training steps, including SFT, RLHF, or even continued pretraining with augmented vocabulary in a new language.
@@ 1 Introduction
Large language models (LLMs) have become the foundation models in many scenarios of artificial intelligence. As training an LLM from scratch consumes a huge amount of computation and data resources and the trained LLM needs to be carefully protected from malicious use, the parameters of the LLMs become a crucial property to protect, for both commercial and ethical reasons. As a result, many of the LLMs are open-sourced with carefully designed licences to reject commercial use (Touvron et al., 2023a; Taylor et al., 2022) or requiring an apply-and-approval process (Touvron et al., 2023b; Zhang et al., 2022; Penedo et al., 2023; BaiChuan-Inc, 2023; Team, 2023; Zheng et al., 2023b), let alone some LLMs are not open-sourced entirely (OpenAI, 2022; GPT-4, 2023; Brown et al., 2020; Wu et al., 2023b; Chowdhery et al., 2022; Hoffmann et al., 2022).
At the core of protecting LLMs from unauthorized use is to identify the base model of a given LLM. However, different from other forms of property such as software or images, protecting LLMs is a novel problem with unique challenges. First, the base model usually needs to be fine-tuned or even continued pretrained to be applied to downstream tasks, resulting in parameter updates that make the resulting model different from the original base model, which makes it disputable to identify the base model. Second, many of the popular LLMs are not releasing their parameters, leaving the identification in a black-box setting. Third, different from previous smaller-scale neural networks that are only trained for specific tasks, LLMs are usually targeted for enormous forms of tasks that are not yet defined during pretraining. This has made the watermarking methods for traditional neural networks (Adi et al., 2018; Xiang et al., 2021; Yadollahi et al., 2021)
not suited in this case, especially under extensive subsequent training steps.
In this work, we propose a novel way to overcome the aforementioned challenges by proposing a model that reads part of the model parameters and computes a fingerprint for each LLM, without directly exposing the LLM parameters to the public or interfering with its training process. The appearance of the fingerprint is closely dependent on the base model, and invariant to almost all subsequent training steps, including supervised fine-tuning (SFT), reinforcement learning with human feedback (RLHF), or even continue-pretraining with augmented vocabulary in a new language.
@@
Moreover, we make the fingerprint human-readable by mapping the invariant terms into a Gaussian random vector through a convolutional encoder and then map the Gaussian vector to a natural image through an off-the-shelf image generation model, StyleGAN2 (Karras et al., 2020). This makes our fingerprints easily interpretable and straightforward to decipher. This is elaborated in Section 4.
@@
Figure 1: An illustrative framework for LLM protection with fingerprints. The LLM manufacturers compute the invariant terms internally, feed them to the public fingerprinting model (FPM) to generate a fingerprint image, and release both to the public. This allows the public to detect shared base models merely from the fingerprint image. Although manufacturers can attempt adjustments on their LLM adversarial to the public FPM to make the fingerprint appear differently, the mere use of random matrices in training the FPM enables the public to train new FPMs whenever necessary, preventing such interference. Importantly, manufacturers only disclose the invariant terms and fingerprint images, without revealing model parameters or affecting LLM training.
@@ 2 Related Works
Despite its short history, safeguarding LLMs against unauthorized use has been a topic of significant interest. There are two primary categories of approaches.
@@
Watermarking Techniques can be divided into two main categories (Boenisch, 2021). The first embeds watermarks or related information into the model parameters, such as explicit watermarking scheme (Uchida et al., 2017) or leveraging another neural network (Wang et al., 2020), which could potentially affect model performance (Wang & Kerschbaum, 2019). The second category focuses on inducing unusual prediction behavior in the model. Xiang et al. (2021) explored embedding phrase triggers, and Gu et al. (2022) extended this approach to LLM; however, they are task-specific. Yadollahi et al. (2021) proposed a watermarking method but did not consider subsequent fine-tuning. Christ et al. (2023) proposed a cryptographic approach, but it is not robust to text editing. Kirchenbauer et al. (2023) involved using pre-selected tokens which inevitably alters the model prediction. These methods may turn out to be vulnerable to attacks on certain tokens, for example, Krishna et al. (2023) successfully evaded watermarking (Kirchenbauer et al., 2023), GPTZero (Tian, 2023), DetectGPT, and OpenAI’s text classifier (OpenAI, 2023) using paraphrasing attacks.
@@ 3.1 Using Vector Direction of LLM Parameters to Identify the Base Model
We can flatten all weight matrices and biases of an LLM into vectors, concatenate all resulting vectors together, and treat is as single huge vector. In this subsection, we are going to show how the direction of this vector could be used to determine the base model by empirically showing its sufficiency and necessity.
@@ 3.1.1 Sufficiency
For sufficiency, we compute the cosine similarities between a base model LLaMA-7B and various of its offspring models, as well as other independently pretrained LLMs (Geng & Liu, 2023) that are of the same size. Table 1 shows a wide spectrum of models that inherit the LLaMA-7B base model, whose subsequent training processes involve various training paradigms, such as SFT (Taori et al., 2023; Xu et al., 2023b; Zheng et al., 2023a; Geng, 2023; Xu et al., 2023a; Han et al., 2023), SFT with LoRA (Wang, 2023) and extensive continued pretraining in a new language (Cui et al., 2023), extending to new modalities (Zhu et al., 2023), etc. We detail the subsequent training settings of these models in the following.
Among these models, Alpaca, Vicuna, Baize, Koala, and WizardLM underwent supervised fine-tuning (SFT) on general datasets of varying sizes and complexity using the LLaMA-7B base model. MedAlpaca performed SFT on medical datasets, while Alpaca-Lora employed the same instruction dataset as Alpaca but used the Lora training method. MiniGPT-4 is a multimodal model that underwent SFT on 5 million aligned image-text pairs. Chinese-Alpaca and Chinese-LLaMA expanded the Chinese vocabulary based on the original LLaMA model, continuing pretraining with 20GB of text data, with Chinese Alpaca further incorporating 2M to 4.3M additional samples during SFT. Additionally, Baichuan, OpenLLaMA, InternLM, and LLaMA2 have identical architectures to LLaMA-7B but were trained from scratch by different organizations or using different training methods.
@@
Model Alpaca-Lora Alpaca Chinese-LLaMA Vicuna Baize MedAlpaca Koala WizardLM MiniGPT-4 Chinese-Alpaca Baichuan OpenLLaMA InternLM LLaMA-2 PCS 99.95 99.98 99.74 99.87 99.80 99.96 99.91 99.96 99.75 99.58 0.95 1.16 1.20 1.51 ICS 99.60 99.95 93.57 99.42 99.60 99.86 99.63 99.89 99.20 91.35 0.32 0.32 0.03 3.16 Table 1: The cosine similarities of model parameters (PCS) and invariant terms (ICS) between various LLMs w.r.t. the LLaMA-7B base model. All models are of the same size.
[Uncaptioned image] Figure 2: The model’s performance quickly deteriorates as the PCS decreases.
@@ 3.1.2 Necessity
From the necessity perspective, we want to verify that if the base model’s ability could still be preserved when the cosine similarity is intentionally suppressed in subsequent training steps. To this end, we inherit the LLaMA-7B base model and interfere with the Alpaca’s SFT process by augmenting the original SFT loss with an extra term that minimizes the absolute value of cosine similarity. i.e. $L_{A}=\frac{\left|\langle\displaystyle{\bm{V}}_{A},\displaystyle{\bm{V}}_{base}\rangle\right|}{|\displaystyle{\bm{V}}_{A}||\displaystyle{\bm{V}}_{base}|}$. Here ${\bm{V}}_{A},\displaystyle{\bm{V}}_{base}$ stand for the parameter vector of the model being tuned and that of the base model, respectively.
Model BoolQ HellaSwag PIQA WinoGrande ARC-e ARC-c RACE MMLU Avg.
LLaMA 75.11 76.19 79.16 70.00 72.90 44.80 40.00 32.75 61.36
Alpaca 77.49 75.64 77.86 67.80 70.66 46.58 43.16 41.13 62.54
$+L_{A}(97.13)$ 45.44 31.16 67.63 48.70 49.03 34.13 22.78 23.13 40.25
$+L_{A}(87.21)$ 42.23 26.09 49.78 47.43 26.43 28.92 22.97 23.22 33.38
$+L_{A}(80.23)$ 39.05 26.40 49.95 48.30 26.52 28.75 22.97 23.98 33.24
$+L_{A}(77.56)$ 41.62 26.15 50.11 49.33 26.56 28.50 22.78 23.12 33.52
Table 2: Detailed zero-shot performance on multiple standard benchmarks of the original LLaMA, Alpaca, and the tuning model at different $L_{A}$(PCS) values.
Figure 2 presents the average zero-shot performance on a set of standard benchmarks when $L_{A}$(PCS) is at different values. The benchmarks include BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018),RACE (Lai et al., 2017) and MMLU (Hendrycks et al., 2020)Table 2 details the performance on each task during SFT. We can see that despite the original training loss is still present, the model quickly deteriorates to random guesses as the cosine similarity detaches away from that of the base model.
@@ 3.2 Deriving the Invariant Terms
Although the vector direction of model parameters is shown to closely stick to its base model, directly comparing the vector direction through cosine similarity requires both models to reveal their parameters, which is unacceptable in many cases. In addition, apart from training, parameter vector direction is vulnerable to some attacks that directly rearrange the model weights. For example, since the hidden units in a model layer are permutation-invariant, one can easily alter the parameter vector direction by randomly permuting the hidden units along with the weights wired to the units. These attacks are invisible to discover since they could easily break the cosine similarity but neither change the model structure nor affect the model performance.
@@
In this subsection, we are going to first systematically analyze and formalize possible weight rearrangements by leveraging the structure constraints of Transformer, and then derive three terms that are invariant under these rearrangements, even when they are combined. Let’s first consider the Transformer layer as depicted in Figure 3. Formally, the layer conducts the following computation:
@@ 3.2.2 The Invariant Terms to These Attacks
In order to find the invariant terms under all these attacks, we need to combine terms in Equation 7 to get the invariant term that nicely cancels out all extra camouflaging matrices. To this end, we construct 3 invariant terms:
@@
Note that we are not simply including all tokens of a vocabulary or tokens in a certain sentence for ${\bm{X}}$, and instead, we use $\hat{\displaystyle{\bm{X}}}$ in these terms. There are two problems if we directly use all tokens’ embeddings ${\bm{X}}$. First, using the whole embedding matrix will make the terms unnecessarily large and of variable size between different models. Second, more importantly, since it is fairly common to inherit a base model with an augmented vocabulary, i.e., to append a set of new tokens at the end of the original vocabulary, the invariant terms will result in different sizes and are not comparable in this case. Third, if we designate specific tokens instead, the selected tokens may not always exist in all LLMs being tested. As a result, we carefully choose the tokens to be included in $\hat{\displaystyle{\bm{X}}}$, by following these steps:
@@
    Tokenize the corpus with the LLM’s vocabulary, and sort all tokens in the vocabulary according to their frequency.
@@
    Among the remaining tokens, select the least frequent $K$ tokens as the tokens to be included in $\hat{\displaystyle\mathbf{X}}$.
Here, using a standard corpus ensures that the resulting tokenization will be identical if a certain model’s vocabulary is a subset of another; the sufficiently large corpus stabilizes the frequencies of tokens in the vocabulary and provides enough chance for as many tokens as possible to show up. Deleting zero-shot tokens automatically sweeps off augmented tokens. Selecting the rarest tokens minimizes potential affections brought by parameter updates in subsequent training processes. A properly large $K$ will ensure a large enough set of tokens are included, making the resulting invariant terms more generally representative. More importantly, it will make all the invariant terms have the same size across all LLMs, regardless of their original sizes. In practice, we choose $K=4096$.
We show the cosine similarity between the invariant terms in Table 1. Although not as perfect as the parameter’s cosine distances, they still preserve a high correlation to the base model.
4 Making The Invariant Terms Human-readable: The Fingerprinting Model
Instead of directly using the three invariant terms, we can present the content in the terms in a human-readable way, through the fingerprinting model. The fingerprinting model consists of a neural network encoder - a convolutional encoder in our case - and an off-the-shelf image generator as depicted in Figure 4. In principle, the encoder takes as input the invariant terms of a certain model, tile them together, and deterministically maps them to a vector that appears to be filled with Gaussian variables. The subsequent image generator reads this vector and maps it to a natural image. Importantly, throughout the process, the locality of the inputs has to be preserved from end to end. i.e., similar invariable terms should result in similar Gaussian variables and finally similar images.

### v3

@@ Abstract
Protecting the copyright of large language models (LLMs) has become crucial due to their resource-intensive training and accompanying carefully designed licenses. However, identifying the original base model of an LLM is challenging due to potential parameter alterations. In this study, we introduce HuRef, a human-readable fingerprint for LLMs that uniquely identifies the base model without interfering with training or exposing model parameters to the public. We first observe that the vector direction of LLM parameters remains stable after the model has converged during pretraining, with negligible perturbations through subsequent training steps, including continued pretraining, supervised fine-tuning, and RLHF, which makes it a sufficient condition to identify the base model. The necessity is validated by continuing to train an LLM with an extra term to drive away the model parameters’ direction and the model becomes damaged. However, this direction is vulnerable to simple attacks like dimension permutation or matrix rotation, which significantly change it without affecting performance. To address this, leveraging the Transformer structure, we systematically analyze potential attacks and define three invariant terms that identify an LLM’s base model. Due to the potential risk of information leakage, we cannot publish invariant terms directly. Instead, we map them to a Gaussian vector using an encoder, then convert it into a natural image using StyleGAN2, and finally publish the image. In our black-box setting, all fingerprinting steps are internally conducted by the LLMs owners. To ensure the published fingerprints are honestly generated, we introduced Zero-Knowledge Proof (ZKP). Experimental results across various LLMs demonstrate the effectiveness of our method. [^1] The code is available at https://github.com/LUMIA-Group/HuRef.
@@ 1 Introduction
Large language models (LLMs) have become the foundation models in many scenarios of artificial intelligence. As training an LLM from scratch consumes a huge amount of computation and data resources and the trained LLM needs to be carefully protected from malicious use, the parameters of the LLMs become a crucial property to protect, for both commercial and ethical reasons. As a result, many of the LLMs are open-sourced with carefully designed licenses to reject commercial use (Touvron et al., 2023a; Taylor et al., 2022) or requiring an apply-and-approval process (Touvron et al., 2023b; Zhang et al., 2022; Penedo et al., 2023; BaiChuan-Inc, 2023; Team, 2023; Zheng et al., 2023b), let alone some LLMs are not open-sourced entirely (OpenAI, 2022; GPT-4, 2023; Brown et al., 2020; Wu et al., 2023b; Chowdhery et al., 2022; Hoffmann et al., 2022).
At the core of protecting LLMs from unauthorized use is to identify the base model of a given LLM. However, different from other forms of property such as software or images, protecting LLMs is a novel problem with unique challenges. First, the base model usually needs to be fine-tuned or even continued pretraining to be applied to downstream tasks, resulting in parameter updates that make the resulting model different from the original base model, which makes it disputable to identify the base model. Second, many of the popular LLMs are not releasing their parameters, leaving the identification in a black-box setting. Third, different from previous smaller-scale neural networks that are only trained for specific tasks, LLMs are usually targeted for enormous forms of tasks that are not yet defined during pretraining. This has made the watermarking methods for traditional neural networks (Adi et al., 2018a; Xiang et al., 2021; Yadollahi et al., 2021) not suited in this case, especially under extensive subsequent training steps.
Figure 1: An illustrative framework for LLM protection with fingerprints. The LLM manufacturers compute invariant terms internally and feed them into the fingerprinting model (FPM [^3] FPM is open-sourced, as all LLM manufacturers need to use the same one. We have placed it on the private side in Figure 1 solely because the fingerprinting process is private.) to generate a fingerprint image. This image is then released to the public along with zero-knowledge proofs ($\pi_{1}$), allowing for intuitive identification of shared base models through the fingerprint images. We also provide a limited one-to-one quantitative comparison scheme (ICS & $\pi_{2}$) as a complement. Zero-Knowledge Proof guarantees the reliability of the fingerprints and comparison results, without interfering with LLM training or revealing model parameters to the public.
In this work, we propose a novel way to overcome the aforementioned challenges by proposing a method that reads part of the model parameters and computes a fingerprint for each LLM without interfering with training or exposing model parameters to the public. The appearance of the fingerprint is closely dependent on the base model, and invariant to almost all subsequent training steps, including supervised fine-tuning (SFT), reinforcement learning with human feedback (RLHF), or even continue-pretraining with augmented vocabulary in a new language.
@@
Moreover, we generate human-readable fingerprints by mapping the invariant terms into a Gaussian random vector through an encoder and then mapping the Gaussian vector to a natural image through an off-the-shelf image generation model, such as StyleGAN2 (Karras et al., 2020). This generation offers a dual benefit of mitigating information leakage and making our fingerprints straightforward to decipher. To ensure the published fingerprints are honestly generated, we also introduced Zero-Knowledge Proof (ZKP). This is elaborated in Section 4.
@@ 2 Related Works
There are two primary categories of related approaches.
@@
Watermarking Techniques can be divided into two main categories (Boenisch, 2021). The first embeds watermarks or related information into the model parameters, such as explicit watermarking scheme (Uchida et al., 2017) or leveraging another neural network (Wang et al., 2020), which could potentially affect model performance (Wang & Kerschbaum, 2019). The second category focuses on inducing unusual prediction behavior in the model. Xiang et al. (2021) explored embedding phrase triggers, and Gu et al. (2022) extended this approach to LLM; however, they are task-specific. Yadollahi et al. (2021) proposed a watermarking method but did not consider subsequent fine-tuning. Christ et al. (2023) proposed a cryptographic approach, but it is not robust to text editing. Kirchenbauer et al. (2023) involved using pre-selected tokens which inevitably alters the model prediction. These methods may turn out to be vulnerable to attacks on certain tokens, for example, Krishna et al. (2023) successfully evaded watermarking (Kirchenbauer et al., 2023), GPTZero (Tian, 2023), and OpenAI’s text classifier (OpenAI, 2023) using paraphrasing attacks.
@@
In Appendix A, we provide a more extensive discussion of additional related works.
Model
Alpaca-Lora
Alpaca
Chinese-LLaMA
Vicuna
Baize
MedAlpaca
Koala
WizardLM
MiniGPT-4
Chinese-Alpaca
Baichuan
OpenLLaMA
InternLM
LLaMA-2
PCS 99.87 99.91 99.68 99.80 99.73 99.90 99.82 99.89 99.70 99.52 0.83 1.16 0.28 1.51
ICS 99.60 99.95 93.57 99.42 99.60 99.86 99.63 99.89 99.20 91.35 0.32 0.32 0.03 3.16
Table 1: The cosine similarities of model parameters (PCS) and invariant terms (ICS) between various LLMs w.r.t. the LLaMA-7B base model. All models are of the same size.
Figure 2: The model’s performance quickly deteriorates as the PCS decreases.
@@ 3.1 Using Vector Direction of LLM Parameters to Identify the Base Model
We can flatten all weight matrices and biases of an LLM into vectors and concatenate them together into a single huge vector. In this subsection, we are going to show how the direction of this vector could be used to determine the base model by empirically showing its sufficiency and necessity.
@@ 3.1.1 Sufficiency
For sufficiency, we compute the cosine similarities between a base model LLaMA-7B and various of its offspring models, as well as other independently pretrained LLMs (Geng & Liu, 2023) that are of the same size. Table 1 shows a wide spectrum of models that inherit the LLaMA-7B base model, whose subsequent training processes involve various training paradigms, such as SFT (Taori et al., 2023; Xu et al., 2023a; Zheng et al., 2023a; Geng, 2023; Xu et al., 2023b; Han et al., 2023), SFT with LoRA (Wang, 2023) and extensive continued pretraining in a new language (Cui et al., 2023), extending to new modalities (Zhu et al., 2023), etc. We detail the subsequent training settings of these models in Appendix Table 10.
@@ 3.1.2 Necessity
From the necessity perspective, we want to verify if the base model’s ability can still be preserved when the cosine similarity is intentionally suppressed in subsequent training steps. To this end, we inherit the LLaMA-7B base model and interfere with the Alpaca’s SFT process by augmenting the original SFT loss with an extra term that minimizes the absolute value of cosine similarity. i.e. $L_{A}=\frac{\left|\langle\displaystyle{\bm{V}}_{A},\displaystyle{\bm{V}}_{base}\rangle\right|}{|\displaystyle{\bm{V}}_{A}||\displaystyle{\bm{V}}_{base}|}$. Here ${\bm{V}}_{A},\displaystyle{\bm{V}}_{base}$ stand for the parameter vector of the model being tuned and that of the base model, respectively.
Figure 2 presents the average zero-shot performance on a set of standard benchmarks when $L_{A}$(PCS) is at different values. The benchmarks include BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), RACE (Lai et al., 2017) and MMLU (Hendrycks et al., 2020). (c.f. Appendix  Table 5 for a detailed breakdown of performances on each task.) We can see that despite the original training loss is still present, the model quickly deteriorates to random guesses as the cosine similarity detaches away from that of the base model.
@@ 3.2 Deriving the Invariant Terms
Although the vector direction of model parameters is shown to closely stick to its base model, directly comparing the vector direction through cosine similarity requires both models to reveal their parameters, which is unacceptable in many cases. In addition, apart from training, parameter vector direction is vulnerable to some attacks that directly rearrange the model weights. For example, since the hidden units in a model layer are permutation-invariant, one can easily alter the parameter vector direction by randomly permuting the hidden units along with the weights wired to the units.
@@
These attacks are invisible to discover since they could easily break the cosine similarity but neither change the model structure nor affect the model performance.
In this subsection, we are going to first systematically analyze and formalize possible weight rearrangements by leveraging the structure constraints of the Transformer, and then derive three terms that are invariant under these rearrangements, even when they are combined. Let’s first consider the Transformer layer as depicted in Figure 3. Formally, the layer conducts the following computation:
@@ 3.2.2 The Invariant Terms to These Attacks
To find the invariant terms under all these attacks, we need to combine terms in Equation 7 to get the invariant term that nicely cancels out all extra camouflaging matrices. To this end, we construct 3 invariant terms:
@@
Note that for $\hat{\displaystyle{\bm{X}}}$ in these terms, we don’t include all tokens from a vocabulary or tokens in a specific sentence; instead, we select a subset of tokens. There are two problems if we directly use all tokens’ embeddings ${\bm{X}}$. First, using the whole embedding matrix will make the terms unnecessarily large and of variable size between different models. Second, more importantly, since it is common to inherit a base model with an augmented vocabulary, i.e., to append a set of new tokens at the end of the original vocabulary, the invariant terms would have different sizes and be incomparable. Third, if we designate specific tokens instead, the selected tokens may not always exist in all LLMs being tested. Consequently, we carefully choose the tokens to be included in $\hat{\displaystyle{\bm{X}}}$, by following these steps:
@@
    Tokenize the corpus with the LLM’s vocabulary and sort all tokens according to their frequency.
@@
    Among the remaining tokens, select the least frequent $K$ tokens as the tokens to be included in $\hat{\displaystyle{\bm{X}}}$.
Here, using a standard corpus ensures that the resulting tokenization will be identical if a certain model’s vocabulary is a subset of another; the sufficiently large corpus stabilizes the frequencies of tokens in the vocabulary and provides enough chance for as many tokens as possible to show up. Deleting zero-shot tokens automatically sweeps off augmented tokens. Selecting the rarest tokens minimizes potential affections brought by parameter updates in subsequent training processes. A properly large $K$ will ensure a large enough set of tokens is included, making the resulting invariant terms more generally representative. More importantly, it will make all the invariant terms have the same size across all LLMs, regardless of their original sizes, i.e., ${\bm{M}}_{a},\displaystyle{\bm{M}}_{b},\displaystyle{\bm{M}}_{f}\in\mathbb{R}^{K\times K}$, regardless of the index of the layer or LLM sizes.
As a result, we can tile up them to form a 3D input tensor ${\bm{M}}\in\mathbb{R}^{K\times K\times C}$, where $C$ is the channel dimension. If we are using all layers, $C=3N$. Again, in order to make ${\bm{M}}$ the same size across all models, we only involve the last $r$ layers in the LLM [^4] In fact, experimentally we find that a small $r$ is already sufficient to discriminate LLMs, it’s not necessary to involve many layers. In all of our experiments, $r=2$, so there are only $6$ channels in the input.. We show the cosine similarity between the invariant terms in Table 1, they still preserve a high correlation to the base model.
4 Mapping the Invariant Terms to Image and Publish it through ZKP
Although invariant terms serve as robust and effective representations for LLMs, we cannot directly publish them due to the potential risk of leaking hidden information, including the size, statistical features, and distribution of parameters. Therefore, we further process invariant terms by mapping them into an image through the fingerprinting model and then publish the image fingerprint instead. This approach helps mitigate the risk of leakage while providing a human-readable fingerprint.
The fingerprinting model consists of a neural network encoder and an off-the-shelf image generator as depicted in Figure 4. In principle, the encoder takes as input the invariant terms of a certain model, tile them together, and deterministically maps them to a vector that appears to be filled with Gaussian variables. The subsequent image generator reads this vector and maps it to a natural image. Importantly, throughout the process, the locality of the inputs has to be preserved from end to end. i.e., similar invariable terms should result in similar Gaussian variables and finally similar images.